The following information for Urbanski Corporation relates to the three months ending June 30, 2018:

Units

Price per unit

Beginning inventory

11,000

$

10

Purchases

75,000

$

16

Sales

80,000

$

25

Ending inventory

6,000

Urbanski uses the LIFO method to account for inventory, and expects at least 15,000 units to be on hand in the ending inventory at year-end. Purchases made in the last six months are expected to cost an average of $18 per unit.

Compute cost of goods sold and gross profit for the quarter ending June 30, 2018.
Answer #1

Urbanski Corporation (LIFO Method)

Cost of goods sold for the quarter ending June 30, 2018, Since it is using LIFO method.

= (75,000 × 16) + (5,000 × 10) (.'. 11,000-6,000 = 5,000)

= 12,00,000 + 50,000

= 12,50,000

Gross profit for the quarter ending June 30,2018: Sales - COGS

= (80,000 × 20) - 12,50,000

= 16,00,000 - 12,50,000

= 3,50,000